The sum of three numbers is 44. The second is three times the first and the third is 6 less than the first. find the numbers.
Found 2 solutions by rapaljer, htmentor:
Answer by rapaljer(4671)

Let x = the first number
3x = second number
x-6= third number

The equation is based upon the SUM of the numbers:
x+ 3x + x-6 = 44
5x - 6 = 44
5x= 50
x=10

x=10 First Number
3x= 3*10=30 Second Number
x-6=4 Third Number

Check: The sum of the numbers is 10 + 30 + 4 = 44 It checks!!

Now, you really need to see my own explanation of WORD PROBLEMS on my own website! To find my website, just click on my tutor name "Rapaljer" anywhere in algebra.com. There you will find a link that takes you right to my Homepage. From my homepage, look for the link "Basic, Intermediate, and College Algebra", then select "Basic Algebra". The word problems are in Chapter 1, so look for Sections 1.09 and 1.10. Most word problems are NOT nearly as hard as they look, if you know how to set them up. This is what I tried to show you with my own "user-friendly" explanations, examples, exercises, ALL with answers too. My own students used to tell me that these are a LOT easier to understand than the traditional textbooks that were written by mathematicians for other mathematicians and publishers! What I have to offer is completely FREE!

Many of the problems in my sections are solved in detail in the corresponding "MATH IN LIVING COLOR" pages. (In fact, the fast way to find my website is to do a Google search for "Math in Living Color", and then take the link at the bottom of this page to go back to my Homepage.)

Anyway, I have a LOT of pages that may be helpful to you including my complete curriculum and a lot of solutions in COLOR to help you understand. There are also VIDEOS (on a variety of topics!) of me teaching in my classes before I retired. Best of all, it's all FREE!

By the way, if you REALLY need this to be solved by Systems of Equations Methods, please send me an Email, and I'll be glad to solve it for you that way.

Dr. R^2

Dr. Robert J. Rapalje
Seminole State College of Florida

Answer by htmentor(1343)

1) x + y + z = 44
2) y = 3x
3) z = x - 6
Substitute the value of y in 2) into equation 1):
x + 3x + z = 44 -> 4x + z = 44
Solve above equation for z in terms of x:
z = 44 - 4x
Substitute the above value for z into equation 3):
44 - 4x = x - 6 -> 5x = 50 -> x = 10
Therefore, from 2), y = 3*10 = 30
And finally, from 3) z = 10 - 6 -> z = 4
Ans: x=10,y=30,z=4